A parabolic reflector has a diameter of 20cm and it's depth is 5cm. Find the focus show solution and graph


1
Expert's answer
2021-09-27T16:10:56-0400


Equation of parabola 

"y^2=2px"

"x=10 => y=\sqrt{2p\cdot 10}=\sqrt{20p}=5=>20p=25=>p=25\20=1.25\\"

Therefore focus F(0,p/2) is F(0,0.625 cm)

Focus lies on the axis of symmetry at a distance 0.625 cm from vertex of parabolic reflector